# Advancing Depth Anything Model for Unsupervised Monocular Depth Estimation in Endoscopy

### v1

@@ IV-E Pose Estimation
In unsupervised monocular depth estimation, camera pose estimation and depth estimation are jointly learned tasks, where the accuracy of the pose network’s estimations can also indirectly reflect the precision of the depth estimation. Therefore, we also evaluated the performance of the pose network. We assessed two video sequences using the Absolute Trajectory Error (ATE) metric, comparing our results only with the most competitive method, EndoDAC [13]. The quantitative results are presented in Table VI, where our method outperforms the compared method across the two sequences. The visual results, shown in Fig. 6, demonstrate reduced trajectory drift and a closer alignment with the ground truth for our method. This further validates the effectiveness of our depth estimation approach.
@@ V CONCLUSIONS
In this work, we apply foundation models to the task of endoscopic image depth estimation. Addressing the challenge of limited medical image data and the inability to train foundation models from scratch, we propose a low-rank adaptation module based on random vectors to fine-tune existing depth estimation foundation model. To overcome the issue of transformers’ weak perception of local information such as edges and textures, we introduce a residual block based on depthwise separable convolution to enhance the network’s capability to capture local features. Comparative experiments show that our method achieves optimal performance on the SCARED dataset with the minimum number of trainable parameters. Ablation studies validate the effectiveness of the proposed modules, and the favorable results from the pose estimation experiments further reflect the accuracy of our approach. Integrating this method into augmented reality navigation systems for minimally invasive endoscopic surgery could significantly enhance the safety and precision of surgical procedures.

### v2

@@ IV-E Pose Estimation
In unsupervised monocular depth estimation, camera pose estimation and depth estimation are jointly learned tasks, where the accuracy of the pose network’s estimations can also indirectly reflect the precision of the depth estimation. Therefore, we also evaluated the performance of the pose network on the SCARED dataset. We assessed three video sequences using the Absolute Trajectory Error (ATE) metric, comparing our results only with the most competitive method, EndoDAC [13]. The quantitative results are presented in Table VII, where our method outperforms the compared method across the three sequences. The visual results, shown in Fig. 7, demonstrate reduced trajectory drift and a closer alignment with the ground truth for our method. This further validates the effectiveness of our depth estimation approach.
@@ V CONCLUSIONS
In this work, we apply foundation models to the task of endoscopic image depth estimation. To address the challenge of limited medical image data and the inability to train foundation models from scratch, we propose a low-rank adaptation module based on random vectors to fine-tune existing depth estimation foundation model. We introduce a residual block based on depthwise separable convolution to enhance the network’s capability to capture local features. Comparative experiments show that our method achieves state-of-the-art performance on the SCARED dataset with the minimum number of trainable parameters. The experimental results on the Hamlyn dataset further validate the strong generalization ability of our method. Ablation studies validate the effectiveness of the proposed modules, and the favorable results from the pose estimation experiments further reflect the accuracy of our approach. Integrating this method into augmented reality navigation systems for minimally invasive endoscopic surgery could enhance surgeons’ spatial awareness of the internal anatomy, thereby improving the safety and precision of the surgical procedure.
Meanwhile, our method has certain limitations in inference time. The large number of parameters in the foundational model results in longer inference times compared to conventional methods. Through testing, our method achieves an average inference time of 30ms per image, with a frame rate of 33 frames per second, which is sufficient to meet real-time requirements. Future work could explore further optimization of the network architecture to achieve more efficient depth estimation.